OpenStudy (anonymous):

Another refresher please. The Midpoint of AC is (-4,16). The coordinates of one endpoint are A(-2,11). Find the coordinates of endpoint C.

jimthompson5910 (jim_thompson5910):

Let C be the point (x,y)

jimthompson5910 (jim_thompson5910):

To get the midpoint of AC, you need to average the corresponding coordinates. So that means we have these two equations (-2+x)/2 = -4 (11+y)/2 = 16 the first equation is for the x coordinate of the midpoint, the second is for the y coordinate of the midpoint

jimthompson5910 (jim_thompson5910):

What do you get when you solve (-2+x)/2 = -4 ?

OpenStudy (anonymous):

i dont understand how to solve that

jimthompson5910 (jim_thompson5910):

(-2+x)/2 = -4 -2+x = -4*2 ... Multiply both sides by 2. -2+x = -8 x - 2 = -8 x = ???

OpenStudy (anonymous):

-6

jimthompson5910 (jim_thompson5910):

yes

jimthompson5910 (jim_thompson5910):

now solve (11+y)/2 = 16

OpenStudy (anonymous):

y=21

jimthompson5910 (jim_thompson5910):

so C is the point (-6, 21)

OpenStudy (anonymous):

thank uu. do uu think you could help me with another geometry question

jimthompson5910 (jim_thompson5910):

sure, I can help with one more

OpenStudy (anonymous):

Find the distance between points A(-5,7) and B(6,-11) to the nearest tenth.

jimthompson5910 (jim_thompson5910):

do you know the distance formula?

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

how far did you get with this formula?

OpenStudy (anonymous):

till the part where d is solved

OpenStudy (anonymous):

i got lost where -11-7 = -18. do i keep the + or keep the -?

jimthompson5910 (jim_thompson5910):

it doesn't matter because when you square -18, you get a positive number

OpenStudy (anonymous):

i got -324 when i squared it though

jimthompson5910 (jim_thompson5910):

you should get +324 not -324

jimthompson5910 (jim_thompson5910):

(-18)^2 = (-18)*(-18) = 324

OpenStudy (anonymous):

say what now

jimthompson5910 (jim_thompson5910):

yeah that's what it means to square a number: you multiply a number by itself

OpenStudy (anonymous):

I knew that but what you had put down confused me. d=17.6

jimthompson5910 (jim_thompson5910):

oh this was confusing? (-18)^2 = (-18)*(-18) = 324

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

which part is confusing?

OpenStudy (anonymous):

just the way it is setup on the computer. Im not use to seeing it like that

jimthompson5910 (jim_thompson5910):

oh probably the ^ symbol huh?

OpenStudy (anonymous):

no i know what that indicates. but anywhooo. i got d=17.5

jimthompson5910 (jim_thompson5910):

17.5 is incorrect

jimthompson5910 (jim_thompson5910):

you should also have 6 - (-5) = 6 + 5 = 11 square that to get 11^2 = 121

jimthompson5910 (jim_thompson5910):

You then add up the squares: 324 + 121 = 445 and take the square root: sqrt( 445 ) = 21.095 that rounds to 21.1

OpenStudy (anonymous):

i had that as my original answer but then i it all in my calculator at one time and yeah it gave me that. but now i see how it isnt

jimthompson5910 (jim_thompson5910):

can you write out how you got 17.5? I'm still not sure how you got that.

OpenStudy (anonymous):

input (6-(-5))^2 + (-11-7)2 and it gets 305 then square root 305 getting 17.46 rounded is 17.5

jimthompson5910 (jim_thompson5910):

odd how it is saying 305 when it should be 445

jimthompson5910 (jim_thompson5910):

(6-(-5))^2 + (-11-7)^2 = 445

OpenStudy (anonymous):

i dont know either,